OpenStudy (anonymous):

using the quadratic formula: x^4-18x^2+17=0

hartnn (hartnn):

replace x^2 by y you will get a quadratic in y

OpenStudy (anonymous):

u=x^2 u^2 =x^4 u^0 = x^0 sub.s in and solve for

OpenStudy (anonymous):

u

OpenStudy (anonymous):

simply put a=1, b= -18 and c= 27 in quadratic formula.. normally quadratic formula is used to determine the value of "x" but here in your case you'll be finding values of "x^2"...

OpenStudy (anonymous):

let y= x^2 y^2=x^4 substitute in there and solve by the formula